# Computing fault-tolerant metric dimension of graphs using their primary subgraphs

### v1

@@ 3 Main results
Since $\operatorname{At}(G_{i})$ is a resolving set for $G_{i}$, for any two vertices of $x_{1},x_{2}\in V(G_{i})$ there exist $a\in\operatorname{At}(G_{i})$, such that $d_{G}(x_{1},a)\neq d_{G}(x_{2},a)$. This implies that there exists a primary end-subgraph $G_{j}$ such that for any $w\in\mathscr{F}_{j}$ we have $d_{G}(a,w)=\min\limits_{v\in V(G_{i})}\{d_{G}(v,w)\}$. Since $G_{j}$ satisfies $\mathscr{C}2$, $|\mathscr{F}_{j}|\geq 2$. Hence, for any $w\in\mathscr{F}_{j}$,
$d_{G}(x_{1},w)=d_{G}(x_{1},a)+d_{G}(a,w_{1})\neq d_{G}(x_{2},a)+d_{G}(a,w)=d_{G}(x_{2},w).$
Case 2: $x_{1}\in V(G_{i})$, $x_{2}\in V(G_{j})$, where $i\neq j$.
Let $a_{1}\in V(G_{i})$ and $a_{2}\in V(G_{j})$ be the attachment vertices such that $d_{G}(x_{1},x_{2})=d_{G}(x_{1},a_{1})+d_{G}(a_{1},a_{2})+d_{G}(a_{2},x_{2}).$ Note that if $G_{i}$ and $G_{j}$ have a common attachment vertex, then $a_{1}=a_{2}$. If $x_{2}=a_{2}=a_{1}$ or $x_{1}=a_{1}=a_{2}$, then we proceed as Case 1. Hence, in the rest we may assume that $x_{1}$ and $x_{2}$ do not belong to the same primary subgraph, that is, $x_{2}\neq a_{1}$ and $x_{1}\neq a_{2}$.
@@
Assume without loss of generality that $|\operatorname{At}(G_{i})|\geq 2$. Since $G_{i}$ satisfies $\mathscr{C}1$, there exists $c\in\operatorname{At}(G_{i})\setminus\{a_{1}\}$, such that $d_{G}(a_{1},c)\geq d_{G}(x_{1},c)$. Now, let $G_{\ell}$, $\ell\neq i$, be a primary end-subgraph such that for any $t\in\mathscr{F}_{\ell}$, $d_{G}(c,t)=\min_{v\in V(G_{i})}\{d_{G}(v,t)\}$, ($\mathscr{F}_{\ell}\geq 2$, as $G_{\ell}$ satisfies $\mathscr{C}2$). Then for any $t_{1},t_{2}\in\mathscr{F}_{\ell}$,
$d_{G}(x_{1},t_{1})$ $=d_{G}(x_{1},c)+d_{G}(c,t_{1})\leq d_{G}(a_{1},c)+d_{G}(c,t_{1})$
$<d_{G}(x_{2},a_{1})+d_{G}(a_{1},c)+d_{G}(c,t_{1})=d_{G}(x_{2},t_{1}),\quad{\rm and}$
$d_{G}(x_{1},t_{2})$ $=d_{G}(x_{1},c)+d_{G}(c,t_{2})\leq d_{G}(a_{1},c)+d_{G}(c,t_{2})$
$<d_{G}(x_{2},a_{1})+d_{G}(a_{1},c)+d_{G}(c,t_{2})=d_{G}(x_{2},t_{2}).$
Subcase 2.2: $|\operatorname{At}(G_{i})|=|\operatorname{At}(G_{j})|=1$.
Clearly $G_{i}$ and $G_{j}$ are primary end-subgraphs and since they satisfy $\mathscr{C}2$, it follows that $|\mathscr{F}_{i}|\geq 2$ and $|\mathscr{F}_{j}|\geq 2$. Hence, let $p_{1},p_{2}\in\mathscr{F}_{i}$ and $q_{1},q_{2}\in\mathscr{F}_{j}$. If there exist two vertices in $\{p_{1},p_{2},q_{1},q_{2}\}$ that distinguish $x_{1}$ and $x_{2}$, then we are done. On the contrary, suppose that there do not exist at least two vertices in $\{p_{1},p_{2},q_{1},q_{2}\}$, such that they individually distinguishes the vertices $x_{1}$ and $x_{2}$. We may suppose without loss of generality that in this case we have:
$d_{G}(x_{1},p_{1})=d_{G}(x_{2},p_{1})=d_{G}(x_{2},a_{2})+d_{G}(a_{2},a_{1})+d_{G}(a_{1},p_{1}),$ (1)
$d_{G}(x_{1},p_{2})=d_{G}(x_{2},p_{2})=d_{G}(x_{2},a_{2})+d_{G}(a_{2},a_{1})+d_{G}(a_{1},p_{2}),$ (2)
$d_{G}(x_{2},q_{1})=d_{G}(x_{1},q_{1})=d_{G}(x_{1},a_{1})+d_{G}(a_{1},a_{2})+d_{G}(a_{2},q_{1}).$ (3)
Observe that since $|\operatorname{At}(G_{i})\cap\operatorname{At}(G_{j})|=\emptyset$, we have $a_{1}\neq a_{2}$. Moreover,
$d_{G}(x_{1},p_{1})\leq d_{G}(x_{1},a_{1})+d_{G}(a_{1},p_{1}),$ (4)
$d_{G}(x_{1},p_{2})\leq d_{G}(x_{1},a_{1})+d_{G}(a_{1},p_{2}),$ (5)
$d_{G}(x_{2},q_{1})\leq d_{G}(x_{2},a_{2})+d_{G}(a_{2},q_{1}).$ (6)
@@
$d_{G}(x_{2},a_{2})+d_{G}(a_{2},a_{1})\leq d_{G}(x_{1},a_{1}).$ (7)
From (3) and (6)
$d_{G}(x_{1},a_{1})+d_{G}(a_{1},a_{2})\leq d_{G}(x_{2},a_{2}).$ (8)
Adding (7) and (8), we get
$2\cdot d_{G}(a_{1},a_{2})\leq 0,$
@@
The following sections focus on deriving several consequences of Theorem 2. Specifically, we present closed-form expressions for the FTMD of certain families of graphs in terms of parameters associated with their primary subgraphs. This is done in cases where the point-attaching process can be described as a graph composition scheme or when the primary subgraphs satisfy specific properties.
@@ 4 An extremal case
As before, let $G$ be a graph formed by point-attaching the primary subgraphs $G_{1},\dots,G_{k}$. In this section, we examine the case where every minimal fault-tolerant basis of a primary subgraph is also of minimum cardinality, that is, when $\operatorname{fdim}(G_{i})=\operatorname{fdim}^{+}(G_{i})$ for each $G_{i}$. Let $\mathscr{F}(G_{i})$ denote the collection of all fault-tolerant metric bases of $G_{i}$ and let
@@
In other words, $\theta_{i}$ represents the maximum number of attachment vertices of $G$ that are also part of a metric basis of $G_{i}$.
@@
Let $G$ be a graph formed by point-attaching $G_{1},\dots,G_{k}$, $k\geq 3$, which satisfy the conditions of Theorem 2. If also $\operatorname{At}(G_{i})\neq V(G_{i})$ and $\operatorname{fdim}(G_{i})=\operatorname{fdim}^{+}(G_{i})$, $i\in[k]$, then
@@
For any primary subgraph $G_{i}$ of $G$ such that $\operatorname{fdim}(G_{i})=\operatorname{fdim}^{+}(G_{i})$, we have $\operatorname{fdim}^{*}(G_{i})=\operatorname{fdim}(G_{i})-\theta_{i}$. Hence, the result follows from Theorem 2. ∎
@@
Figure 2: A graph $G$ formed by point-attaching from $G_{1}\cong K_{4}$, $G_{2}\cong K_{3}$, $G_{3}$ being the paw graph, $G_{4}\cong C_{8}$, and $G_{5}\cong K_{5}$.
A block graph is a graph in which each biconnected component (block) forms a clique. Observe that any block graph can be constructed by point-attaching a family of connected graphs. Recalling that $\operatorname{fdim}(K_{n})=n=\operatorname{fdim}^{+}(K_{n})$, we get the following remark as a special case of Corollary 1.
@@
Let $G$ be a block graph formed by point-attaching $\{K_{r_{1}},K_{r_{2}},\dots,K_{r_{k}}\}$, where $k\geq 3$ and $r_{i}\geq 3$, $i\in[k]$. If $\operatorname{At}(K_{r_{i}})\cap\operatorname{At}(K_{r_{j}})=\emptyset$ for any primary end-subgraphs $K_{r_{i}}$ and $K_{r_{j}}$, then,
@@ 5 Rooted products
In this section, we explore a notable special case of graphs constructed through point-attaching: the rooted product of graphs.
A rooted graph is a graph with a designated vertex that is uniquely labeled to distinguish it from the others. This distinguished vertex is referred to as the root of the graph. Let $G$ be a labeled graph with $n$ vertices, and let $\mathscr{H}=\{H_{1},\dots,H_{n}\}$ be a collection of rooted graphs. The rooted product graph $G[\mathscr{H}]$ is formed by identifying the root of each graph $H_{i}$ with the $i^{\text{th}}$ vertex of $G$. It is evident that any rooted product graph $G[\mathscr{H}]$ can be viewed as a graph obtained by point-attaching. Here, the primary internal subgraph is $G$ with $\operatorname{At}(G)=V(G)$, while the family $\mathscr{H}$ consists of primary end-subgraphs, where each attachment vertex corresponds to the root of the respective graph. Using Theorem 2, we derive the following result.
@@
Let $G$ be a connected graph of order $n\geq 2$, and let $\mathscr{H}=\{H_{1},\dots,H_{n}\}$ be a family of rooted graphs, each satisfying $\mathscr{C}2$, with roots $v_{1},\dots,v_{n}$, respectively. Then
@@
where $H_{i}\in\mathscr{H}_{1}$ if $v_{i}$ is not part of any fault-tolerant basis of $H_{i}$, and $H_{j}\in\mathscr{H}_{2}$ otherwise.
Next, we examine the case where the family $\mathscr{H}$ consists of vertex-transitive graphs. Let $\text{Aut}(H)$ denote the automorphism group of a graph $H$. For any two vertices $x_{1},x_{2}\in V(H)$ and any automorphism $f\in\text{Aut}(H)$, the distance between the vertices is preserved, i.e., $d(x_{1},x_{2})=d(f(x_{1}),f(x_{2}))$. Consequently, if $\mathscr{F}$ is a fault-tolerant basis of a connected graph $H$ and $f\in\text{Aut}(H)$, then the image of the basis under the automorphism, $f(\mathscr{F})$, is also a fault-tolerant basis of $H$. Thus, each vertex of $H$ must belongs to some fault-tolerant basis. Applying Corollary 1, we derive the following remark.
@@
Let $\mathscr{H}=\{H_{1},\dots,H_{n}\}$ be a collection of vertex-transitive graphs with orders greater than two. Then for any connected graph $G$ of order $n\geq 2$, we have
@@
In particular, if $\mathscr{H}=\{K_{r_{1}},\dots,K_{r_{n}}\}$, then
@@
A special case of rooted product graphs arises when the family $\mathscr{H}$ consists of $n$ isomorphic rooted graphs. Formally, let $V(G)=\{u_{1},\dots,u_{n}\}$, and let $v$ be the root vertex of a graph $H$. The rooted product graph $G\circ_{v}H$ is defined with the vertex set $V(G\circ_{v}H)=V(G)\times V(H)$ and the edge set
$E(G\circ_{v}H)=\bigcup_{i=1}^{n}\{(u_{i},b)(u_{i},y):by\in E(H)\}\cup\{(u_{i},v)(u_{j},v):u_{i}u_{j}\in E(G)\}.$
For this case, Corollary 1 reduces to the following.
@@
If $H$ is a connected graph not isomorphic to a path and $v\in V(H)$, then the following hold.
@@
    If $v$ is not a part of any fault-tolerant basis of $H$, then for any connected graph $G$ of order $n$,
@@
    If $v$ belongs to a fault-tolerant basis of $H$, then for any connected graph $G$ of order $n\geq 2$,
@@
Proposition 3 raises the question of identifying the necessary and/or sufficient conditions for a vertex $v\in V(H)$ to be part of a fault-tolerant basis of $H$. For example, it is straightforward to verify that a vertex $v$ in a path graph $P$ belongs to fault-tolerant basis of $P$ if and only if $v$ is one of its leaf vertices. Building on this observation, Proposition 3 (i) yields:
@@
Let $H$ be a connected graph, $v\in V(H)$ a vertex that is not part of any fault-tolerant basis of $H$, and let $G$ be a connected graph of order $n$. Then $\operatorname{fdim}(G\circ_{v}H)=2n$ if and only if $H$ is a path graph and $v$ is not a leaf.
@@
If $G$ is a connected graph of order $n\geq 2$, and $v$ is a leaf of a non-trivial path $P$, then
@@
$G$ appears as an induced subgraph of $G\circ_{v}P$. Since any fault-tolerant resolving set of $G\circ_{v}P$ must in particular resolve $G$, the lower bound follows.
To establish the upper bound we claim that $V(G)\times\{v^{\prime}\}$ is a fault-tolerant resolving set for $G\circ_{v}P$, where $v^{\prime}$ denotes the leaf of $P$ distinct from $v$. Let $(x,y)$ and $(x^{\prime},y^{\prime})$ be any two vertices of $G\circ_{v}P$. If $x=x^{\prime}$, then for any $u_{1},u_{2}\in V(G)$,
@@
Hence $V(G)\times\{v^{\prime}\}$ is indeed a fault-tolerant resolving set for $G\circ_{v}P$, hence $\operatorname{fdim}(G\circ_{v}P)\leq n$. ∎

### v2

@@ 3 Main results
Since $\operatorname{At}(G_{i})$ is a resolving set for $G_{i}$, for any two vertices of $x_{1},x_{2}\in V(G_{i})$ there exist $a\in\operatorname{At}(G_{i})$, such that $d(x_{1},a)\neq d(x_{2},a)$. This implies that there exists a $G_{j}$ such that for each $w\in\mathscr{F}_{j}$ we have $d(w,a)=\min\limits_{v\in V(G_{i})}\{d(w,v)\}$. Now, $|\mathscr{F}_{j}|\geq 2$ due to $G_{j}$ satisfying $\mathscr{C}2$. Hence, for any $w\in\mathscr{F}_{j}$,
$d(x_{1},w)=d(x_{1},a)+d(a,w_{1})\neq d(x_{2},a)+d(a,w)=d(x_{2},w).$
Case 2: $x_{1}\in V(G_{i})$, $x_{2}\in V(G_{j})$, $G_{i}\neq G_{j}$.
Let $a_{1}\in\operatorname{At}(G_{i})$ and $a_{2}\in\operatorname{At}(G_{j})$, such that $d(x_{1},x_{2})=d(x_{1},a_{1})+d(a_{1},a_{2})+d(a_{2},x_{2}).$ Note that $a_{1}=a_{2}$, whenever $\operatorname{At}(G_{i})\cap\operatorname{At}(G_{j})\neq\emptyset$. If $x_{2}=a_{2}=a_{1}$ or $x_{1}=a_{1}=a_{2}$, then the discussion is similar to Case 1. Hence, in the rest we may assume that $x_{2}\neq a_{1}$ and $x_{1}\neq a_{2}$.
@@
Let $|\operatorname{At}(G_{i})|>1$. As $G_{i}$ obeys $\mathscr{C}1$, $\exists\ c\in\operatorname{At}(G_{i})\setminus\{a_{1}\}$, such that $d_{G}(a_{1},c)\geq d_{G}(x_{1},c)$. Assume $G_{\ell\neq i}$, with $|\operatorname{At}(G_{\ell\neq i})|=1$, s.t. for every $t\in\mathscr{F}_{\ell}$, $d(t,c)=\min_{v\in V(G_{i})}\{d(t,v)\}$, ($\mathscr{F}_{\ell}\geq 2$, as $G_{\ell}$ obeys $\mathscr{C}2$). Then for every $t_{1},t_{2}\in\mathscr{F}_{\ell}$,
$d(x_{1},t_{1})$ $=d(c,x_{1})+d(c,t_{1})\leq d(c,a_{1})+d(c,t_{1})$
$<d(x_{2},a_{1})+d(c,a_{1})+d(c,t_{1})=d(x_{2},t_{1}),\quad{\rm and}$
$d(x_{1},t_{2})$ $=d(c,x_{1})+d(c,t_{2})\leq d(c,a_{1})+d(c,t_{2})$
$<d(x_{2},a_{1})+d(c,a_{1})+d(c,t_{2})=d(x_{2},t_{2}).$
Subcase 2.2: $|\operatorname{At}(G_{i})|=1=|\operatorname{At}(G_{j})|$.
Since, $G_{i}$ and $G_{j}$ obeys $\mathscr{C}2$, $|\mathscr{F}_{i}|\geq 2$ and $|\mathscr{F}_{j}|\geq 2$. Hence, let $p_{1},p_{2}\in\mathscr{F}_{i}$ and $q_{1},q_{2}\in\mathscr{F}_{j}$. If there exist two vertices in $\{p_{1},q_{1},p_{2},q_{2}\}$ that distinguish $x_{1}$ and $x_{2}$, then we are done. Suppose that there do not exist at least two vertices in $\{p_{1},q_{1},p_{2},q_{2}\}$, s.t. they individually distinguishes the vertices $x_{1}$ and $x_{2}$. We may suppose w.l.o.g. that in this case we have:
$d(x_{1},p_{1})=d(x_{2},a_{2})+d(a_{2},a_{1})+d(a_{1},p_{1})=d(x_{2},p_{1}),$ (1)
$d(x_{1},p_{2})=d(x_{2},a_{2})+d(a_{2},a_{1})+d(a_{1},p_{2})=d(x_{2},p_{2}),$ (2)
$d(x_{2},q_{1})=d(x_{1},a_{1})+d(a_{1},a_{2})+d(a_{2},q_{1})=d(x_{1},q_{1}).$ (3)
Observe that since $\operatorname{At}(G_{i})\cap\operatorname{At}(G_{j})=\emptyset$, we have $a_{1}\neq a_{2}$. Moreover,
$d(x_{1},p_{1})\leq d(p_{1},a_{1})+d(a_{1},x_{1}),$ (4)
$d(x_{1},p_{2})\leq d(p_{2},a_{1})+d(a_{1},x_{1}),$ (5)
$d(x_{2},q_{1})\leq d(q_{1},a_{2})+d(a_{2},x_{2}).$ (6)
@@
$d(a_{1},a_{2})+d(a_{2},x_{2})\leq d(x_{1},a_{1}).$ (7)
From (3) &\& (6)
$d(x_{1},a_{1})+d(a_{2},a_{1})\leq d(x_{2},a_{2}).$ (8)
Adding (7) &\& (8), we get
$2\cdot d(a_{1},a_{2})\leq 0,$
@@
Let us now illustrate Theorem 3 with two examples, the first of which is a block graph.
Example 4.
The graph $G$ illustrated in Figure 4 satisfies conditions $(\mathscr{C}_{1})$ and $(\mathscr{C}_{2})$. Hence, Theorem 3 can be applied to compute the fault-tolerant metric dimension of $G$ using the attaching metric dimensions of its primary subgraphs as follows:
$\operatorname{fdim}(G)$ $=\operatorname{fdim}^{*}(G_{1})+\operatorname{fdim}^{*}(G_{2})+\operatorname{fdim}^{*}(G_{3})+\operatorname{fdim}^{*}(G_{4})+\operatorname{fdim}^{*}(G_{5})$
$=2+4+3+2+4$
[…]
Figure 4: A block graph $G$ constructed by point attaching $G_{1}\cong K_{3}$, $G_{2}\cong K_{6}$, $G_{3}\cong K_{4}$, $G_{4}\cong K_{3}$, and $G_{5}\cong K_{5}$.
Example 5.
The graph $G$ illustrated in Figure 5 satisfies conditions $(\mathscr{C}_{1})$ and $(\mathscr{C}_{2})$. Hence, Theorem 3 can be applied to compute the fault-tolerant metric dimension of $G$ using the attaching metric dimensions of its primary subgraphs as follows:
$\operatorname{fdim}(G)$ $=\operatorname{fdim}^{*}(G_{1})+\operatorname{fdim}^{*}(G_{2})+\operatorname{fdim}^{*}(G_{3})+\operatorname{fdim}^{*}(G_{4})+\operatorname{fdim}^{*}(G_{5})$
$=2+4+0+2+2$
[…]
Figure 5: A graph $G$ constructed by point attaching $G_{1}\cong C_{7}$, $G_{2}\cong K_{6}$, $G_{3}\cong P_{3}$, $G_{4}\cong C_{4}$, and $G_{5}\cong K_{3}$.
The following sections focus on deriving several consequences of Theorem 3. Specifically, we present closed-form expressions for the FTMD of certain collection of graphs which are derived by point-attaching process.
@@ 4 An extremal case
The investigation of the case where each minimal fault-tolerant basis of a primary subgraph is also of minimum cardinality, that is, when $\operatorname{fdim}(G_{i})=\operatorname{fdim}^{+}(G_{i})$ for each $G_{i}$ were done in this section. Let $\mathscr{F}(G_{i})$ denote the collection of all FTB of $G_{i}$ and let
@@
Let $G$ is formed by the point-attaching process over $\{G_{i}:i\in[k]\}$, $k\geq 3$, which satisfy the conditions of Theorem 3. If also $\operatorname{At}(G_{i})\neq V(G_{i})$ and $\operatorname{fdim}(G_{i})=\operatorname{fdim}^{+}(G_{i})$, $i\in[k]$, then
@@
By the assumption, for any $G_{i}$ we have $\operatorname{fdim}(G_{i})=\operatorname{fdim}^{+}(G_{i})$. Note now that the identity $\operatorname{fdim}^{*}(G_{i})=\operatorname{fdim}(G_{i})-\theta_{i}$ follows by the definition $\theta_{i}=\max_{F\in\mathscr{F}(G_{i})}|F\cap\operatorname{At}(G_{i})|$, because for any attaching fault-tolerant resolving set of $G_{i}$, the remaining vertices must necessarily be chosen from $V(G_{i})\setminus\operatorname{At}(G_{i})$, and their number is therefore $\operatorname{fdim}(G_{i})-\theta_{i}$. Hence, the result follows from Theorem 3. ∎
@@
Figure 6: A graph $G$ and its primary subgraphs $G_{1}\cong K_{4}$, $G_{2}\cong K_{3}$, $G_{3}$ being the paw graph, $G_{4}\cong C_{8}$, and $G_{5}\cong K_{5}$.
It is interesting to see that block graph can be derived by point-attaching process on a collection of complete graphs. Recalling that $\operatorname{fdim}(K_{n})=n=\operatorname{fdim}^{+}(K_{n})$, we get the remark as a special case of Corollary 1, which is as follows.
@@
If $G$ is a block graph formed by the point-attaching process over $\{K_{r_{i}}:i\in[k]\}$, where $k\geq 3$ and $r_{i}\geq 3$. If $\operatorname{At}(K_{r_{i}})\cap\operatorname{At}(K_{r_{j}})=\emptyset$ for any primary end-subgraphs $K_{r_{i}}$ and $K_{r_{j}}$, then,
@@ 5 Rooted products
In this section, we explore a rooted product of graphs constructed through point-attaching.
A rooted graph is the one with a designated vertex that is uniquely vertex labeled to distinguish it from the others. This distinguished vertex is referred to as the root of the graph. Let $G$ be a vertex labeled graph of order $n$, and let $\mathscr{H}=\{H_{i}:i\in[n]\}$ be a collection of rooted graphs. The rooted product graph denoted by $G[\mathscr{H}]$ is formed by identifying the root of each $H_{i}$ with the $i^{\text{th}}$ vertex of $G$ as defined in [44]. It is evident that any rooted product graph $G[\mathscr{H}]$ can be viewed as a graph derived by point-attaching process. Using Theorem 3, we derive the following result.
@@
Let $G$ of order $n\geq 2$, and let $\mathscr{H}=\{H_{i}:i\in[n]\}$ be a collection of rooted graphs, each satisfying $\mathscr{C}2$, with roots $v_{1},\dots,v_{n}$, respectively. Then
@@
where $H_{i}\in\mathscr{H}_{1}$ if $v_{i}$ is not part of any FTB of $H_{i}$, and $H_{j}\in\mathscr{H}_{2}$ otherwise.
Next, we examine the case where the family $\mathscr{H}$ consists of graphs that are vertex-transitive. Let $\text{Aut}(H)$ denote the automorphism group of a graph $H$. For any two vertices $x_{1},x_{2}\in V(H)$ and any automorphism $f\in\text{Aut}(H)$, the distance between the vertices is preserved. Consequently, if $\mathscr{F}$ is a FTB of $H$ and $f\in\text{Aut}(H)$, then the image of the basis under the automorphism, $f(\mathscr{F})$, is also a FTB of $H$. Thus, each vertex of $H$ must belong to some FTB. The next remark is the direct consequence of Corollary 1.
@@
Let $\mathscr{H}=\{H_{i}:i\in[n]\}$ be a collection of vertex-transitive graphs of order at least 3, and $G$ of order at least 2, then
@@
Further, if $\mathscr{H}=\{K_{r_{i}}:i\in[n]\}$, we have
@@
A special case of rooted product graphs arises when the family $\mathscr{H}$ consists of $n$ isomorphic rooted graphs. Let $V(G)=\{g_{i}:i\in[n]\}$, and let $V(H)=\{h_{i}:i\in[n^{\prime}]\}$. Declare the vertex $h=h_{1}$ to be the root of $H$. The rooted product $G\circ_{h}H$ has the vertex set
$V(G\circ_{h}H)=V(G)\times V(H)=\big\{(g_{i},h_{j}):\ i\in[n],j\in[n^{\prime}]\big\}\,$,
and the edge set
$E(G\circ_{h}H)=\big\{(g_{i},h)(g_{i^{\prime}},h):\ g_{i}g_{i^{\prime}}\in E(G)\big\}\cup\bigcup_{i=1}^{n}\big\{(g_{i},h_{j})(g_{i},h_{j^{\prime}}):\ h_{j}h_{j^{\prime}}\in E(H)\big\}\,.$
The following result is a special case of Corollary 1.
@@
If $H$ is not isomorphic to a path and $v\in V(H)$, then the following hold.
@@
    If $v$ is not a part of any FTB of $H$, then for any $G$ of order $n$,
@@
    If $v$ belongs to a FTB of $H$, then for any $G$ of order at least 2,
@@
Proposition 6 raises the question of identifying the conditions of necessity and sufficiency for $v\in V(H)$ to be part of a FTB of $H$. It is straightforward to verify that a $v$ is in FTB of $P$ iff $v$ is one of its leaf vertices. Building on this observation, Proposition 6 (i) yields:
@@
Let $v\in V(H)$ is not part of any FTB of $H$, and let $G$ be a graph of order $n$. Then $\operatorname{fdim}(G\circ_{v}H)=2n$ iff $H\cong P$ and $v$ is not a leaf.
@@
If $G$ is of order $n\geq 2$, and $v$ is a leaf of a non-trivial path $P$, then
@@
$G$ appears as an induced subgraph of $G\circ_{v}P$. Since any FTRS of $G\circ_{v}P$ must in particular resolve $G$, the lower bound follows.
To establish the upper bound we claim that $V(G)\times\{v^{\prime}\}$ is a FTRS of $G\circ_{v}P$, $v^{\prime}\neq v$ denotes the leaf of $P$. Let $(x^{\prime},y^{\prime}),(x,y)\in V(G\circ_{v}P)$. If $x=x^{\prime}$, then for any $u_{1},u_{2}\in V(G)$,
@@
Hence $V(G)\times\{v^{\prime}\}$ is indeed a FTRS for $G\circ_{v}P$, hence $\operatorname{fdim}(G\circ_{v}P)\leq n$. ∎